OpenStudy (anonymous):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 33 feet and the child ride has a radius of 22 feet, what is the ratio between the volumes of each ride?

OpenStudy (anonymous):

33/22 = 3/2 ... sooo since volume ratio=(radius ratio)^3 (3/2)^3 = 7 7/8 hmm.... im not 100% sure is this one of your options?

OpenStudy (anonymous):

or well i guess 63/8 as a improper fraction